May 27, 2014 BCAL Commissioning

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:15 a.m. Eastern

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


Action Items

  1. None.

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Updates
  4. Any other business


Attendees: Elton, Will M., Laura, Shayne (JLab), Andrei, Tegan, Zisis (Regina)

  1. Announcements: The HallD staff has moved to the A-wing. Elton's office is A108.
  2. Updates
    1. Shayne and Laura produced a new schematic connecting the BCAL geometry (sectors, etc) to the fADC crates, channels. This has been pasted next to the crates in the Hall for easy locating of channels.
    2. Will: a cosmics run was taken over this past weekend (106K events). Andrei's bottom counters were used in coincidence with the top counter to form the trigger. See logbook entry for more details.
    3. Hardware fixes (south side): Nick and Chris have repaired all problem channels. There was a variety of issues: a) HV#1 was not getting to the module: connector was damaged and fixed. b) set voltage was not being reached: faulty MPOD module which was replaced. c) module with dead channel: pinched cables that were fixed.
    4. Software/controls: LED and bias controlled was locked; Hovannes will be contacted.
  3. Upcoming tasks
    1. Check fixes with LED pulser data on the south as well as the north side.
    2. Keep monitoring pedestals, periodic fluctuations with temperatures in hall/crates.
    3. Check ADV values of trigger scintillators to ensure their HV is set correctly.
    4. Take more cosmics runs when everything is working.
      1. Check energy depositions in cells, starting with vertically oriented ones.
      2. Check gains, including individual SiPMs into the sums.
    5. Develop scripts for analysis (students, Andrei)
  4. APD project: Elton and students used the Hall B facility (it test up to 32 APDs at a time) to test 4 APDs. The system includes a chiller and temperature studies can be made. A calibrated LED is used. IV curves are extracted. The APDs are contenders for the rare eta decay calorimeter.
  5. Adjourned.